# Unifying uncertainties for rotor-like quantum systems

### 04 Mar 2024 v1

@@ Appendix A Derivation of moment of inertia for arbitrary axis
$M=M_{0}-|\langle C\rangle|^{2}\cos^{2}\theta.$ (50)
Using the relation $D^{2}=1-|\langle C\rangle|^{2}$ and $D^{2}=\langle[\Delta S(\Phi)]^{2}\rangle+\langle[\Delta C(\Phi)]^{2}\rangle$ we get finally the expression where the moment of inertia is given as weighted sum of both cosine and sine variances
$M=D^{2}-\langle[\Delta C(\Phi)]^{2}\rangle\sin^{2}\theta$ (51)
$=\langle[\Delta S(\Phi)]^{2}\rangle+\langle[\Delta C(\Phi)]^{2}\rangle\cos^{2}\theta.$ (52)
@@ Appendix B The tight uncertainty relations for generic moment of inertia
There are two ways motivated by previous approach. There are three parameters of the covariance matrix (13), e.g., its trace, determinant and relative phase between moments or just moments $\langle E\rangle,\langle E^{2}\rangle$ (apart of an overall arbitrary phase) which can be exploited for quantification of the ”noise” associated with the covariance matrix.
The first formulation is motivated by a similar approach as in (6) seeking the minimum of the variance of angular momentum under the constraint of fixed values of moments $\langle E\rangle$ and $\langle E^{2}\rangle$
@@
$\mu,q,r$ being Lagrange multipliers. Since we are seeking the solution for $\langle L\rangle=0$ and zero phase, we might set $\mu=0$ an in $\phi$-representation this tends to Hill equation, similarly to previously discussed problem of Mathieu function,
@@
$\beta=\arg[\langle E^{2}\rangle\langle E\rangle^{*2}].$ Inequality (53) should be further modified in a similar way as it was done in (17) in order to identify the angular uncertainty term.
The second formulation will require to minimize the product $\langle(\Delta L)^{2}\rangle M_{\bf n}$ for fixed values of $\theta,\Phi.$ This corresponds to fixed values $D^{2}$ and $\langle(\Delta C(\Phi))^{2}\rangle,$ it means just two constraints in Hill equation. Finally, the product should be minimised with respect to $\theta,\Phi,$ which should be finally identified as state dependent quantities
$\langle(\Delta L)^{2}\rangle M_{\bf n}\geq C(\langle E\rangle,\langle E^{2}\rangle).$ (56)
Notice by passing that von Mises state - a solution of Eq. (12), can be cast as a special case of Hill equation corresponding to the ground state of the Hamiltonian
$\left(\Delta L+i\kappa\Delta S_{\alpha}\right)|\left(\Delta L-i\kappa\Delta S_{\alpha}\right)|\psi\rangle=0.$ (57)
@@ C.2 Bounds for uncertainty product $\Pi_{\mathcal{S}}$
By the numerical minimization of the uncertainty product (59) over the parameter $\kappa_{s}$ at fixed $\kappa_{a}$ we get the black line in Fig. 4 of the main text. The line again possesses a sharp point but now for a slightly different value of $\kappa_{a,{\rm sh}}\doteq 2.897$, for which $\tilde{D}_{a,{\rm sh}}^{2}\doteq 0.174$. Comparison with the numerically minimized uncertainty product (59) reveals that in the interval $1\geq D_{a}^{2}\geq\tilde{D}_{a,{\rm sh}}^{2}$ the second branch $\langle(\Delta L_{a})^{2}\rangle/2$ is equal to the minimal uncertainty product, whereas for $\tilde{D}_{a,{\rm sh}}^{2}>D_{a}^{2}\geq 0$ the first branch lies slightly above the minimal uncertainty product (c.f. black line and magenta line in Fig. 4 of the main text).

### 13 Mar 2024 v2

@@ Appendix A Elementary derivation of moment of inertia for arbitrary axis
$M=M_{0}-|\langle C\rangle|^{2}\cos^{2}(\theta).$ (50)
Using the relation $D^{2}=1-|\langle C\rangle|^{2}$ and $D^{2}=\langle(\Delta S_{\Phi})^{2}\rangle+\langle(\Delta C_{\Phi})^{2}\rangle$ we get finally the expression where the moment of inertia is given as weighted sum of both cosine and sine variances
$M=D^{2}-\langle(\Delta C_{\Phi})^{2}\rangle\sin^{2}(\theta)$ (51)
$=\langle(\Delta S_{\Phi})^{2}\rangle+\langle(\Delta C_{\Phi})^{2}\rangle\cos^{2}(\theta).$ (52)
Notice that measures discussed in Sec. II are extremal -dispersion is the largest whereas $\langle(\Delta S_{\Phi})^{2}\rangle$ is the smallest.
@@ Appendix B The tight uncertainty relations for generic moment of inertia
There are three parameters of the covariance matrix (13), e.g., its trace, determinant and relative phase between moments or just moments $\langle E\rangle,\langle E^{2}\rangle$ (apart of an overall arbitrary phase) which can be exploited for quantification of the ”noise” associated with the covariance matrix.
The first formulation is motivated by a similar approach as in inequality (6) seeking the minimum of the variance of angular momentum under the constraint of fixed values of moments $\langle E\rangle$ and $\langle E^{2}\rangle$
@@
$\mu,q,r$ being Lagrange multipliers. Since we are seeking the solution for $\langle L\rangle=0$ and zero phase, we might set $\mu=0$ an in $\phi$-representation this tends to Hill equation [27], similarly to previously discussed problem of Mathieu function,
@@
$\beta=\arg[\langle E^{2}\rangle\langle E\rangle^{*2}].$ Inequality (53) does not have the form of uncertainty relations since it does not contain angular uncertainty term.
Uncertainty relations based on the moment of inertia can be considered as minimum of the product $\langle(\Delta L)^{2}\rangle M_{\bf n}$ for the fixed values of $\theta,\Phi$ under the constraint of fixed $D^{2}$ and $\langle(\Delta C_{\Phi})^{2}\rangle.$ Extremal states are again solutions of Hill equation. Finally, the product should be minimised with respect to $\theta,\Phi$ and these parameters should be identified as state dependent quantities analogously to construction adopted in derivation of inequalities (17) saturated by von Mises states.
Notice that von Mises state - a solution of Eq. (12), can be also cast as a special case of (factorized) Hill equation corresponding to the ground state of the Hamiltonian
$\left(\Delta L+i\kappa\Delta S_{\alpha}\right)\left(\Delta L-i\kappa\Delta S_{\alpha}\right)|\psi\rangle=0.$ (56)
@@ C.2 Bounds for uncertainty product $\Pi_{\mathcal{S}}$
By the numerical minimization of the uncertainty product (58) over the parameter $\kappa_{s}$ at fixed $\kappa_{a}$ we get the black line in Fig. 4 of the main text. The line again possesses a sharp point but now for a slightly different value of $\kappa_{a,{\rm sh}}\doteq 2.897$, for which $\tilde{D}_{a,{\rm sh}}^{2}\doteq 0.174$. Comparison with the numerically minimized uncertainty product (58) reveals that in the interval $1\geq D_{a}^{2}\geq\tilde{D}_{a,{\rm sh}}^{2}$ the second branch $\langle(\Delta L_{a})^{2}\rangle/2$ is equal to the minimal uncertainty product, whereas for $\tilde{D}_{a,{\rm sh}}^{2}>D_{a}^{2}\geq 0$ the first branch lies slightly above the minimal uncertainty product (c.f. black line and magenta line in the white area of Fig. 4 of the main text).